Determine the concentration of CO2 in water at 15 degrees C when the concentration of CO2 in the atmosphere is 0.02%
Answer #1

Apply Henry's Law first...

Concentration in mol/L = KH​ x p​i

where KH ​is the Henry constant and pi is the mole fraction of CO2 in air

Get Pi, patial pressure of CO2

P-CO2 = mol frac of CO2 * Ptotal

Assume Ptotal = 1 atm

so the mol frac --> 0.02/100 = 0.0002

so,

P-CO2 = 0.0002*1 = 0.0002 atm

M​= (3.5*10^-2) ​* ( 0.0002) =0.000007 mol of CO2/L